OpenStudy (greengalaxy):

Simplify the expression. -4 + 6 x -1 A. –10 B. –2 C. 2 D. 10 Now someone out there helped me understand this but im checking this awnser, I think its c:2

OpenStudy (greengalaxy):

@Awolflover1

OpenStudy (greengalaxy):

@hartnn

OpenStudy (onepieceftw):

-4+6 is just 6-4 and multiplying by -1 makes it negative.

OpenStudy (greengalaxy):

So -4+6 = 2 and 2 x -1 = -2 '

OpenStudy (onepieceftw):

correct

OpenStudy (kaylardz0405):

right

OpenStudy (greengalaxy):

Ah thank you, so it is C?

OpenStudy (onepieceftw):

It's B -2

OpenStudy (greengalaxy):

Oh B? So that means i left out the -4.

OpenStudy (onepieceftw):

-4+6=2 2 times -1 = -2 o the answer is B, -2

OpenStudy (greengalaxy):

Oh thank you very much! I gave you a medal. going to close this question

OpenStudy (onepieceftw):

YW